Monday Homework Problem 1.4 Your standard number 2 mechanical pencil has a graphite lead with mass 0.05g. How many protons are in this quantity of graphite? The atomic mass of carbon is 12g/mole and the atomic number is 6. Select One of the Following: (a) 3.1 x 1010 protons (b) 1.5 x 1022 protons (c) 5.1 x 1024protons (d) 2.3 x 1021 protons (e) 6.5 x 1023 protons Monday Homework Problem 1.5 The dielectric plate of an electrophorus is charged by rubbing with an insulating material. The neutral conducting plate of the electrophorus is then held by its insulating handle and placed on the charged dielectric. The metal plate remains neutral through this process. Select the choice that follows that best describes the distribution of charge on the conducting plate of the electrophorus as it sits on the charged dielectric plate. Select One of the Following: (a) There is no net density charge anywhere on the electrophorus. (b) There is a uniform surface charge on the electrophorus with the same sign as the charge of the dielectric (c) There is a uniform volume charge spread throughout the electrophorus with the same sign as the charge of (d) There is a surface charge opposite the sign of the charge of the dielectric plate on the side of the conducting plate. the dielectric plate. plate in contact with the dielectric, and a surface charge with the same sign as that of the dielectric on the other side. (e) There is a uniform surface charge on the electrophorus with the opposite sign as the charge of the dielectric plate.
Monday Homework Problem 1.7 Two neutral conducting spheres, S and S2. are connected by a conducting wire. A positively charged rod is brought near Si but does not transfer charge to S1. The wire is cut and then the rod removed. After the rod has been removed, what is the sign of the charge on S2? Select One of the Following (a) The charge of S2 is approximately zero. (b) positive (c) negative
Answer #1

1.4)

m = given mass = 0.05 g

M = atomic mass = 12 g/mole

n = number of moles

number of moles are given as

n = m/M

n = 0.05/12 = 0.00417

No = Avagadro's number = 6.023 x 1023

N = number of atoms = n No = (0.00417) ( 6.023 x 1023) = 2.5 x 1021

Ne = total number of protons = 6 N

Ne = 6 (2.5 x 1021)

Ne = 1.5 x 1022 protons

b)